10 years ago, Al was half as old as Bill is now, and 7 years from now, Al will be as old as Bill is now. Set up a system of equations that can be used to determine the ages now of Al and Bill, then solve the system
Answer by henew(13)

Al's current age A
Bill's current age B.
"10 years ago, Al was half as old as Bill is now" A-10= B/2

"7 years from now, Al will be as old as Bill is now" A+7=B
B/2=A-10 can be written as B= 2A-20 ( multiply by 2 on both sides) . From second equation B= A+7
2A-20= A+7
2A-A=20+7
SO
A=27
B= 27+7=34